Find the roots of the following quadratic equations by factorisation: 
 2x ^2 + x – 6 = 0 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear student

Given,
2x^2 + x – 6 = 0
Taking LHS,
=> 2x^2 + 4x - 3x - 6
=> 2x(x + 2) - 3(x + 2)
=> (x + 2)(2x - 3)
The roots of this equation, 2x^2 + x – 6=0
are the values of x for which (x - 5)(x + 2) = 0
Therefore, x + 2 = 0 or 2x - 3 = 0
=> x = -2 or x = 3/2

Thanks